OpenStudy (anonymous):

solve the equation if possible. check ur answers. |x+4|=3 someone help me!!

OpenStudy (espex):

If you move the abs to the other side you will get x+4=|3|, so \[x+4=\pm 3\]

OpenStudy (espex):

Solve for x in both cases.

OpenStudy (anonymous):

do you need help for further steps?

OpenStudy (anonymous):

|x+4|=3 | < means it will always be positive so yeah subtract 4 from each side: x= -1 that's how I understand it! :>

OpenStudy (xanthe):

...

OpenStudy (zzr0ck3r):

|x+4| = 3 --> x+4 = 3 and x+4 = -3 so x = -1 and x = -7